OpenStudy (anonymous):

What formula do i use for this? How many vertices does a convex polyhedron have if it has 14 faces and 24 edges?

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

This is my first time here.

OpenStudy (mayaal):

F+V=E-2

OpenStudy (anonymous):

Ok but i doesnt give me the vertices though?

OpenStudy (mayaal):

F+faces V=vertices E=edges

OpenStudy (anonymous):

So face will be 14 and edges will be 24, but what do i put for vertices

OpenStudy (anonymous):

I did, then i got 36 but my paper says its either 8, 10, 11 ,12

OpenStudy (anonymous):

As answers.

OpenStudy (anonymous):

Im comfused on this one?!

OpenStudy (mayaal):

Look 14+24=E-2

OpenStudy (anonymous):

Ok

OpenStudy (mayaal):

Solve for E instead of V sorry i accidentally told you to solve for V

OpenStudy (mayaal):

Now 14+24 is 38

OpenStudy (anonymous):

Its 40=E

OpenStudy (anonymous):

Correct?

OpenStudy (mayaal):

Then it will be 38=E-2

OpenStudy (anonymous):

I get this part but my paper says the answers may be 8, 10, 11, a12

OpenStudy (mayaal):

We will have to do something with it...

OpenStudy (anonymous):

*12

OpenStudy (anonymous):

But i get 40

OpenStudy (mayaal):

Lets use this formula:F + V - E = 2

OpenStudy (anonymous):

So I am really stuck here!

OpenStudy (anonymous):

Like what?

OpenStudy (mayaal):

Yup it is 40

OpenStudy (anonymous):

Is there another formula?

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

What do i put for v

OpenStudy (anonymous):

And it cant be 40 if my options is 8, 10, 11,12

OpenStudy (mayaal):

Nothing just plug in the ones you are given

OpenStudy (anonymous):

I am only given 2 Edges and faces but i dont have the vertices

OpenStudy (mayaal):

F + V - E = 2 would be 14+V-24=2

OpenStudy (anonymous):

And i plug them in ill get 40 but 40 is not a option. the options it gives me is 8,10,11,12

OpenStudy (anonymous):

OHHHHHHHHHHHHHHH

OpenStudy (mayaal):

Got it...?

OpenStudy (anonymous):

Brain fart!

OpenStudy (anonymous):

Is it 8?

OpenStudy (mayaal):

-10+V=2

OpenStudy (mayaal):

V=12

OpenStudy (mayaal):

so it is 12.....

OpenStudy (anonymous):

It would be -8

OpenStudy (mayaal):

:) :) now did u get it?

OpenStudy (anonymous):

cause -10+2=-8

OpenStudy (anonymous):

Oh you did! -10+V=2 +10 +10 V=12

OpenStudy (mayaal):

No, you will add 10 to both sides so that 10 will be gone from the side where there is -10 because -10+10 is 0...

OpenStudy (anonymous):

Ohh you so helpful :D

OpenStudy (mayaal):

Yeah thats what I did

OpenStudy (anonymous):

Ok can you list the names of the 2 formulas you did?

OpenStudy (anonymous):

So i can write it down?

OpenStudy (mayaal):

F + V - E = 2

OpenStudy (mayaal):

I dont think the other one will work....

OpenStudy (anonymous):

Ok the First one is to solve the one with no Vertices

OpenStudy (anonymous):

And this one F+V=E-2 will solve if i have all the info for F, V AND E

OpenStudy (anonymous):

Thanks agaain i gave you gold medal ;D

OpenStudy (mayaal):

let me seee

OpenStudy (mayaal):

thanks so much

OpenStudy (mayaal):

wait let me see the other formula

OpenStudy (anonymous):

Np you deserve i might be posting again only IF I GET SUCK AGAIN. Dont want to abuse this wonderful system :D

OpenStudy (anonymous):

Stuck*

OpenStudy (anonymous):

Ill wait :D

OpenStudy (anonymous):

6 vertices, 8 faces, and 12 edges I think this formula can solve this type. F+V=E-2

OpenStudy (mayaal):

yeah that one will work tooo......! F+V=E-2

OpenStudy (anonymous):

:D

OpenStudy (mayaal):

but i would rather use the one i posted before.

OpenStudy (anonymous):

What would the answer be for this one?

OpenStudy (mayaal):

if u need my help just post @mayaal where you post ur answers...

OpenStudy (mayaal):

the answer is 12

OpenStudy (anonymous):

Ok thanks again!

OpenStudy (anonymous):

Will do!

OpenStudy (mayaal):

r u in algebra?

OpenStudy (anonymous):

Geometry

OpenStudy (mayaal):

where do you live?

OpenStudy (mayaal):

i am in geometry too...!

OpenStudy (anonymous):

Oooh i live in CA

OpenStudy (mayaal):

I live in Florida

OpenStudy (anonymous):

Lol

OpenStudy (mayaal):

lol..........! :)

OpenStudy (anonymous):

I got to unversal studios hollywood you go to orlandos ;P

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

Well bye!

OpenStudy (mayaal):

YEAH

OpenStudy (anonymous):

XD

OpenStudy (mayaal):

BYE BYE